OpenStudy (anonymous):

Becky and Tommy are asked to solve -4x - 25 = 8x + 15. Identify where one of them made an error. (1 point) Becky Tommy -4x - 25 = 8x + 15 -4x - 25 - 8x = 8x + 15 - 8x -12x - 25 = 15 -12x - 25 + 25 = 15 + 25 -12x = 40 40 40 -3/10 -4x - 25 = 8x + 15 -4x - 25 + 4x = 8x + 15 + 4x -25 = 12x + 15 -25 - 15 = 12x + 15 - 15 -40 = 12x 12 12 -10/3 Question 2 options: 1) Becky made the error when she subtracted 8x 2) Becky made the error when she divided by 40 3) Tommy made the error when he added 4x 4) Tommy made the error when he added 15

OpenStudy (texaschic101):

Do you have any thoughts on this one ? Hint : take a close look at Becky's work

DivineSolar (divinesolar):

Ah becky made the mistake when she divided 40 as i have solved the problem 2. Let's solve your equation step-by-step. −4x−25=8x+15 Step 1: Subtract 8x from both sides. −4x−25−8x=8x+15−8x −12x−25=15 Step 2: Add 25 to both sides. −12x−25+25=15+25 −12x=40 Step 3: Divide both sides by -12. −12x/−12=40/−12 x=−10/3 Answer: x=−10/3

DivineSolar (divinesolar):

Refresh the page.